# Data-informativity conditions for structured linear systems with implications for dynamic networks

### v2

@@ 7.2 Discussion
The new single module data-informativity results presented in Theorem 8 and Proposition 9, are relaxed versions of the data-informativity conditions presented in Van den Hof et al. (2023), where the target was to consistently estimate the full (unstructured) predictor model. It has been shown in Van den Hof et al. (2023) that for satisfying those data-informativity conditions it may be necessary to adjust the selection of predictor inputs and outputs. In other words, for a particular selection of predictor inputs and outputs, it is not guaranteed that the data-informativity can be met by adding a sufficient number of external excitation signals. That is different in the current results, where the data-informativity conditions can always be satisfied by adding a sufficient number of external excitation signals at appropriate locations.
@@
Consider the situation of Theorem 8. Then for any predictor model the data-informativity condition for consistent estimation of the target module $\bar{G}_{ji}^{0}(q)$ can be satisfied, by adding a sufficient number of external excitation signals to the predictor inputs $w_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}$.
@@
When estimating a single module $G_{ji}^{0}(q)$ in a dynamic network, the user has the option now to require two different sets of conditions for data-informativity, either the single module conditions if the current paper, or the full predictor model conditions of Van den Hof et al. (2023). While the single module conditions are more relaxed, this situation also comes at a cost. When validating an estimated module, as part of a consistently estimated predictor model, model validation can be applied to the full predictor model, for validating all of the present modules. This can e.g. be done by considering a validation data set. However when only the target module is estimated consistently, validation of this module using a validation data set, would require that in this validation data set, the node signals that are input to parametrized modules in the predictor model, are exactly the same as in the estimation data set. This is caused by the fact that in the consistent single module estimation case, the estimated modules, different from the target module, will become dependent on the particular input node signals, present in the estimation data set.
@@
So far we have been dealing with the data-informativity conditions for the direct method of estimating predictor models in a dynamic network, see Ramaswamy and Van den Hof (2021). There are alternative methods, as e.g. the multistage method of Fonken et al. (2023), where data-informativity conditions are also more relaxed. Construction of predictor models for both the direct method and the multistage method, including the different types of data-informativity conditions, has been implemented in the MATLAB app and toolbox SYSDYNET Van den Hof et al. (2024).
@@
One of the prime consequences of the results presented in this paper is in the relation with single module identifiability. Identifiability conditions for single modules in a dynamic network have been studied in Weerts et al. (2018a); Hendrickx et al. (2019); Weerts et al. (2018b); Shi et al. (2022, 2023). These conditions are typically not dependent on a particular identification algorithm. At the same time, identification algorithms have been studied that provide consistent target module estimates under specific conditions, among which data-informativity conditions. One would expect that if the single module identifiability conditions are satisfied, that there exists an estimation algorithm that can provide the consistent target module estimate. However, so far this step has only been achieved for a particular situation, namely when the single module identifiability conditions are satisfied by external excitation signals $r$ only. In that situation an indirect method can be used to estimate the target module consistently, see e.g. Van den Hof et al. (2013); Dankers et al. (2016); Gevers et al. (2018). However when both $r$-excitation and external $e$-signals are required for satisfying the single module identifiability conditions, there was a gap with the related data-informativity conditions of the estimation algorithms.
@@
In order to understand how the current results close this gap, we consider the identifiability result as formulated in Theorem 4 and Corollary 1 of Shi et al. (2022), adapted here to the situation of single module identifiability. We also consider a model set $\mathcal{M}$ of a full network that satisfies the basic assumptions as formulated in Shi et al. (2022), including the assumption that all modules in $G(q,\theta)$ are strictly proper.
Theorem 11 (Shi et al. (2022))
Let $\mathcal{X}_{j}$ be composed of those components in $x=(u,e)$ that in the model set $\mathcal{M}$ do not have a parametrized link to $w_{j}$. Then $G_{ji}$ is generically identifiable in $\mathcal{M}$ from $(w,r)$ if there exists a disconnecting set $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ from $\{\mathcal{X}_{j}\cup w_{i}\}$ to $w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}\backslash\{w_{i}\}$, subject to $i\notin\mathcal{D}^{c}$, such that there exist $dim(w_{\!{\scriptscriptstyle\mathcal{D}}^{c}})+1$ vertex disjoint paths from $\mathcal{X}_{j}$ to $w_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}$.
Now it appears that the conditions for generic single module data-informativity, as formulated in Proposition 9 are extremely closely related to the generic single module identifiability conditions of Theorem 11. While in Proposition 9 external signals in set $x_{\!{\scriptscriptstyle\mathcal{T}}}$ are excluded from excitation, this similar situation is covered in Theorem 11 by requiring that the disconnecting set $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ should also be a disconnecting set from $\mathcal{X}_{j}$ to $w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}\backslash\{w_{i}\}$. I.e. external signals that reach $w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}\backslash\{w_{i}\}$ without passing through $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ are excluded. This set is exactly covered by $w_{\!{\scriptscriptstyle\mathcal{T}}}$.
@@
The single difference between the conditions in the two results, is in terms of the white noise signals $e$ that are available for excitation. While in Theorem 11 only white noise signals with a parametrized link to $w_{j}$ are excluded, in Proposition 9 these are the white noise signals with a parametrized path to $w_{\!{\scriptscriptstyle\mathcal{Y}}}$. In the situation of a single output (MISO) predictor model these sets are exactly the same. However in the multi-output case, there is apparently some conservatism in the direct identification method. Choosing a multiple output predictor model for identification of a single module is typically done for coping with confounding variables, see e.g. Ramaswamy and Van den Hof (2021). Increasing the number of outputs apparently comes at the cost of more strict data-informativity conditions. This is in line with the data-informativity conditions for consistency of full predictor models, as presented in Ramaswamy and Van den Hof (2021); Van den Hof et al. (2023). Alternative methods, as the multi-stage method (Fonken et al. (2023)) do not seem to suffer from this conservatism.
@@ 8.1 Two-node example
We consider the 2-node example as depicted in Figure 3 (Van den Hof and Ramaswamy, 2020) with target module $G_{21}$.
@@
Because of the confounding variables that occur due to the correlated disturbances, we include node signal $w_{1}$ in the output, leading to a predictor model $w_{1}\rightarrow(w_{1},w_{2})$. In the original reasoning of the local direct method, data-informativity is analyzed for the full SIMO model, leading to the situation that neither $u_{1}$ nor $u_{2}$ can be used for excitation of input $w_{1}$. This is due to the fact that in modelling output $w_{1}$ in the predictor model, and since $w_{2}$ is not an input signal, both $u_{1}$ and $u_{2}$ appear in the predictor expression for $w_{1}$ with unknown dynamic terms, reflecting the sensitivity of the feedback loop between $w_{1}$ and $w_{2}$. The solution was then to move to a 2-input 2-output model $(w_{1},w_{2})\rightarrow(w_{1},w_{2})$ which then satisfies the data-informativity conditions when both $u_{1}$ and $u_{2}$ are present Van den Hof et al. (2023). In the analysis of the current paper, we consider the predictor model $w_{1}\rightarrow(w_{1},w_{2})$, but for data-informativity only need to account for the mapping $w_{1}\rightarrow w_{2}$. In this case $w_{i}=w_{1}$, $w_{j}=w_{2}$, $\mathcal{T}=\emptyset$ and $\mathcal{D}^{c}=\emptyset$. Since in the predictor model for predicting $w_{2}$, $u_{1}$ is not present, and $u_{2}$ appears with a constant term $1$, it follows that $u^{[j]}=\emptyset$, and therefore both $u_{1}$ and $u_{2}$ can be used to excite $w_{1}$. As a result, data-informativity is guaranteed generically if either $u_{1}$ or $u_{2}$ is present. This is in agreement with the earlier observations in Van den Hof et al. (2017).
@@ 8.2 Six-node example
We consider the 6-node example as depicted in Figure 4 with target module $G_{12}$.
@@
Given that the target module is $G_{12}$ we start composing the predictor model with input $w_{2}$ and output $w_{1}$. Because of the parallel path connection between $w_{2}$ and $w_{1}$ through $w_{3}$, we add $w_{3}$ as predictor input. And in order to handle the confounding variable that is present through the correlated disturbances $v_{1},v_{2}$, we add $w_{1}$ to the predictor output. This leads to a predictor model $(w_{2},w_{3})\rightarrow(w_{1},w_{2})$, that satisfies the structural conditions for consistently estimating the target module. With this choice, $w_{i}=w_{2}$ and $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}=w_{3}$.
When analyzing the data-informativity conditions as developed in this paper, the vector signal $(w_{2},w_{3})$ would need appropriate excitation. Let us first consider possible locations $(x_{5},x_{3},x_{6})$ for unmeasured disturbance signals. Possible disturbance signals $x_{4}=e_{4}$ and $x_{6}=e_{6}$ will not provide excitation for $(w_{2},w_{3})$ as these disturbance signals have unmeasured paths to one of the output signals $w_{2}$ or $w_{1}$. Therefore their white noise sources are part of $e^{\!{\scriptscriptstyle\mathcal{Y}}}$, and cannot be used for excitation. A disturbance signal $x_{3}=e_{3}$ will lead to excitation for the input signal $w_{3}$.
@@
If it would not be possible to add an external signal to $w_{3}$, one could wonder whether that could be replaced by an external signal added to $w_{6}$. As mentioned above, an unmeasured disturbance signal $e_{6}$ would not be exciting, as it would have an unknown path to the output $w_{1}$, and therefore needs to be excluded from excitation. Similarly, an external excitation signal $u_{6}$ would have an unknown path output $w_{1}$, and therefore would be in $u^{\perp 1}$, and thus be excluded from exciting the inputs $(w_{2},w_{3})$. The conclusion therefore is that in the considered situation, an external signal added to $w_{3}$ is required.
When applying the earlier results for full predictor model data-informativity, as presented in Van den Hof et al. (2023), the result would be that the present predictor model cannot satisfy the data-informativity conditions by adding a sufficient number of external signals only. Relying on the result of Proposition 5 in Van den Hof et al. (2023), we would need to extend the predictor model to $(w_{1},w_{2},w_{3},w_{5})\rightarrow(w_{1},w_{2})$, thus adding $w_{5}$ as input node signal to be measured, and also adding $w_{1}$ to the predictor input. This is based on the result of the mentioned Proposition, indicating that each loop in teh network around nodes that appear both in the input and the output, should pass through an input node. Sufficient conditions for (full predictor model) data-informativity would then be that $w_{5}$ and $w_{3}$ are excited by either a noise signal $(e_{3},e_{5})$ or an excitation signals $(u_{3},u_{5})$, and that both $w_{1}$ and $w_{2}$ are excited by an external excitations signals $u_{1}$ and $u_{2}$. This illustrates how single module data-informativity can lead to significantly more relaxed conditions for instrumenting the network with sensors and actuators.
@@ 9 Conclusions
The components of $x$ that appear in the equation for $u_{k}$, $k\in\{i\cup\mathcal{D}^{c}\cup\mathcal{T}\}$ are then the $x$ components that are directly added to the concerned notes in the original system representation, plus all those components $x$ that only pass through nodes in $u_{\!{\scriptscriptstyle\mathcal{F}}}$ before they reach $u_{k}$. This is a direct result of Gaussian elimination (immersion) of the nodes $u_{\!{\scriptscriptstyle\mathcal{F}}}$ (Dankers et al., 2016).
@@
Since according to (17), $u_{\!{\scriptscriptstyle\mathcal{T}}}$ can be written as a function of $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ and $x_{\!{\scriptscriptstyle\mathcal{T}}}$, substituting this in the above expression provides
@@
$\cdot\begin{bmatrix}u_{i}\\ u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\\ x_{\!{\scriptscriptstyle\mathcal{T}}}\end{bmatrix}.$
Denote $\nu:=\begin{bmatrix}u_{i}\\ u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\end{bmatrix}$. Then we can decompose $\nu$ according to $\nu=\nu_{x_{\!{\scriptscriptstyle\mathcal{T}}}}+\nu_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}}}$ and using the fact that $\Delta_{j}^{b}(q,\theta)$ is parametrized independently from $\Delta G_{ji}$ and $\Delta_{j}^{a}$, we can rewrite $\Delta\bar{M}_{j}^{u}(q,\theta)\kappa_{u}^{[j]}$ as
@@
$\cdot\begin{bmatrix}\nu_{x_{\!{\scriptscriptstyle\mathcal{T}}}}+\nu_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}}}\\ x_{\!{\scriptscriptstyle\mathcal{T}}}\end{bmatrix}$
The expression for $\Delta\bar{M}_{j}^{u}(q,\theta)\kappa_{u}^{[j]}$ is now written as a summation of two terms that are mutually uncorrelated: one driven by $x_{\!{\scriptscriptstyle\mathcal{T}}}$ and $\nu_{x_{\!{\scriptscriptstyle\mathcal{T}}}}$ and one driven by its orthogonal complement. According to (38), for evaluating $\Delta\bar{M}_{j}(q,\theta)\kappa^{[j]}(t)$ a third term is added, given by $\Delta\bar{M}_{j}^{e}(q,\theta)\kappa_{e}^{[j]}$, leading to
@@
$\left(\begin{bmatrix}v_{x_{\!{\scriptscriptstyle\mathcal{T}}}}\\ x_{\!{\scriptscriptstyle\mathcal{T}}}\end{bmatrix}+\begin{bmatrix}v_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}}}\\ 0\end{bmatrix}\right)+\Delta\bar{M}_{j}^{e}(q,\theta)\kappa_{e}^{[j]}.$
Since both terms $v_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}}}$ and $\kappa_{e}^{[j]}$ are orthogonal to $x_{\!{\scriptscriptstyle\mathcal{T}}}$, it follows now that $\bar{\mathbb{E}}\|\Delta\bar{M}_{j}(q,\theta)\kappa^{[j]}(t)\|^{2}=0$ implies $\begin{bmatrix}\Delta G_{ji}(q,\theta)&\Delta_{j}^{c}(q,\theta)&\Delta\bar{M}_{j}^{e}(q,\theta)\end{bmatrix}=0$ under the condition that $\Phi_{\mu^{[j]}}(\omega)\succ 0$ for almost all $\omega$ with $\mu^{[j]}$ given by
$\mu^{[j]}=\begin{bmatrix}\nu_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}}}\\ \kappa_{e}^{[j]}\end{bmatrix},$
@@
$\Delta\bar{M}_{j}^{w}(q,\theta)=\begin{bmatrix}\Delta\bar{G}_{ji}(q,\theta)&\Delta_{j}^{a}(q,\theta)&\Delta_{j}^{b}(q,\theta)\end{bmatrix}$ (42)
as
$\Delta\bar{M}_{j}^{u}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}=\begin{bmatrix}\Delta\bar{G}_{ji}(q,\theta)&\Delta_{j}^{a}(q,\theta)&\Delta_{j}^{b}(q,\theta)\end{bmatrix}\begin{bmatrix}w_{i}\\ w_{\bar{\!}{\scriptscriptstyle\mathcal{D}}^{c}}\\ w_{\!{\scriptscriptstyle\mathcal{T}}}\end{bmatrix}$ (43)
with $w_{\bar{\!}{\scriptscriptstyle\mathcal{D}}^{c}}=w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\cap w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}$. By complementing $w_{\bar{\!}{\scriptscriptstyle\mathcal{D}}^{c}}$ with the remaining components of $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ and adding 00 entries to the accompanying $\Delta^{a}_{j}(q,\theta)$, this can be rewritten as
$\Delta\bar{M}_{j}^{u}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}=\begin{bmatrix}\Delta\bar{G}_{ji}(q,\theta)&\tilde{\Delta}_{j}^{a}(q,\theta)&\Delta_{j}^{b}(q,\theta)\end{bmatrix}\begin{bmatrix}w_{i}\\ w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\\ w_{\!{\scriptscriptstyle\mathcal{T}}}\end{bmatrix}.$ (44)
Since according to (32), $w_{\!{\scriptscriptstyle\mathcal{T}}}$ can be written as a function of $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ and $x_{\!{\scriptscriptstyle\mathcal{T}}}$, substituting this in the above expression provides

### v3

@@ 8.2 Discussion
The new single module data-informativity results presented in Theorem 8 and Proposition 9, are relaxed versions of the data-informativity conditions presented in VandenHof&etal_IFAC:23 , where the target was to consistently estimate the full (unstructured) predictor model. It has been shown in VandenHof&etal_IFAC:23 that for satisfying those data-informativity conditions it may be necessary to adjust the selection of predictor inputs and outputs. In other words, for a particular selection of predictor inputs and outputs, it is not guaranteed that the data-informativity can be met by adding a sufficient number of external excitation signals. That is different in the current results, where the data-informativity conditions can always be satisfied by adding a sufficient number of external excitation signals at appropriate locations.
@@
Consider the situation of Theorem 8. Then for any predictor model that satisfies the structural conditions for module invariance and confounding variables handling [^7] See conditions a. and b. in Section 8.1., there exists a disconnecting set $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ such that the data-informativity condition for consistent estimation of the target module $\bar{G}_{ji}^{0}(q)$ can be satisfied, by adding a sufficient number of external excitation signals to the predictor inputs $w_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}$.
@@
When estimating a single module $G_{ji}^{0}(q)$ in a dynamic network, the user has the option now to require two different sets of conditions for data-informativity, either the single module conditions of the current paper, or the full predictor model conditions of VandenHof&etal_IFAC:23 . While the single module conditions are more relaxed, this situation also comes at a cost. When validating an estimated module, as part of a consistently estimated predictor model, model validation can be applied to the full predictor model, for validating all of the present modules. This can e.g. be done by considering a validation data set. However when only the target module is estimated consistently, validation of this module using a validation data set, would require that in this validation data set, the node signals that are input to parametrized modules in the predictor model, are exactly the same as in the estimation data set. This is caused by the fact that in the consistent single module estimation case, the estimated modules, different from the target module, will become dependent on the particular input node signals, present in the estimation data set.
@@
So far we have been dealing with the data-informativity conditions for the direct method of estimating predictor models in a dynamic network, see Ramaswamy&VandenHof_TAC:21 . There are alternative methods, as e.g. the multi-step method of Fonken&etal_CDC:23 , where data-informativity conditions are also more relaxed. Construction of predictor models for both the direct method and the multi-step method, including the different types of data-informativity conditions, has been implemented in the MATLAB app and toolbox SYSDYNET VandenHof&etal_SYSID:24 .
@@
One of the prime consequences of the results presented in this paper is in the relation with single module identifiability. Identifiability conditions for single modules in a dynamic network have been studied in Weerts&etal_Autom:18_identifiability ; Hendrickx&Gevers&Bazanella_TAC:19 ; Shi&etal_Autom:22 ; Shi&etal_TAC:23 . These conditions are typically not dependent on a particular identification algorithm. At the same time, identification algorithms have been studied that provide consistent target module estimates under specific conditions, among which data-informativity conditions. One would expect that if the single module identifiability conditions are satisfied, that there exists an estimation algorithm that can provide the consistent target module estimate. However, so far this step has only been achieved for a particular situation, namely when the single module identifiability conditions are satisfied by external excitation signals $r$ only. In that situation an indirect method can be used to estimate the target module consistently, see e.g. VandenHof&etal_Autom:13 ; Dankers&etal_TAC:16 ; Gevers&etal_sysid:18 . However when both $r$-excitation and external $e$-signals are required for satisfying the single module identifiability conditions, there was a gap with the related data-informativity conditions of the estimation algorithms.
@@
In order to understand how the current results close this gap, we consider the identifiability result as formulated in Theorem 4 and Corollary 1 of Shi&etal_Autom:22 , adapted here to the situation of single module identifiability. We also consider a model set $\mathcal{M}$ of a full network that satisfies the basic assumptions as formulated in Shi&etal_Autom:22 , including the assumption that all modules in $G(q,\theta)$ are strictly proper.
Theorem 11 (Shi&etal_Autom:22 )
[…]
@@
The single difference between the conditions in the two results, is in terms of the white noise signals $e$ that are available for excitation. While in Theorem 11 only white noise signals with a parametrized link to $w_{j}$ are excluded, in Proposition 9 these are the white noise signals with a parametrized path to $w_{\!{\scriptscriptstyle\mathcal{Y}}}$. In the situation of a single output (MISO) predictor model these sets are exactly the same. However in the multi-output case, there is apparently some conservatism in the direct identification method. Choosing a multiple output predictor model for identification of a single module is typically done for coping with confounding variables, see e.g. Ramaswamy&VandenHof_TAC:21 . Increasing the number of outputs apparently comes at the cost of more strict data-informativity conditions. This is in line with the data-informativity conditions for consistency of full predictor models, as presented in Ramaswamy&VandenHof_TAC:21 ; VandenHof&etal_IFAC:23 . The multi-step method (Fonken&etal_CDC:23 ) does not seem to suffer from this conservatism, due to an alternative handling of confounding variables.
@@ 9.1 Two-node example
We consider the 2-node example as depicted in Figure 4 (VandenHof&Ramaswamy_CDC:20, ) with target module $G_{21}$.
@@
Because of the confounding variables that occur due to the correlated disturbances, we include node signal $w_{1}$ in the output, leading to a predictor model $w_{1}\rightarrow(w_{1},w_{2})$. In the original reasoning of the local direct method, data-informativity is analyzed for the full SIMO model, leading to the situation that neither $u_{1}$ nor $u_{2}$ can be used for excitation of input $w_{1}$. This is due to the fact that in modelling output $w_{1}$ in the predictor model, and since $w_{2}$ is not an input signal, both $u_{1}$ and $u_{2}$ appear in the predictor expression for $w_{1}$ with unknown dynamic terms, reflecting the sensitivity of the feedback loop between $w_{1}$ and $w_{2}$. The solution was then to move to a 2-input 2-output model $(w_{1},w_{2})\rightarrow(w_{1},w_{2})$ which then satisfies the data-informativity conditions when both $u_{1}$ and $u_{2}$ are present VandenHof&etal_IFAC:23 . In the analysis of the current paper, we consider the predictor model $w_{1}\rightarrow(w_{1},w_{2})$, but for data-informativity only need to account for the mapping $w_{1}\rightarrow w_{2}$. In this case $w_{i}=w_{1}$, $w_{j}=w_{2}$, $\mathcal{T}=\emptyset$ and $\mathcal{D}^{c}=\emptyset$. Since in the predictor model for predicting $w_{2}$, $u_{1}$ is not present, and $u_{2}$ appears with a constant term $1$, it follows that $u^{[j]}=\emptyset$, and therefore both $u_{1}$ and $u_{2}$ can be used to excite $w_{1}$. As a result, data-informativity is guaranteed generically if either $u_{1}$ or $u_{2}$ is present.
@@ 9.2 Six-node example
We consider the 6-node example with target module $G_{21}$, presented in Section 2, while for ease of reference we repeat its figure as Figure 5.
@@
Given that the target module is $G_{12}$ we start composing the predictor model with input $w_{2}$ and output $w_{1}$. Because of the parallel path connection between $w_{2}$ and $w_{1}$ through $w_{3}$, we add $w_{3}$ as predictor input. And in order to handle the confounding variable that is present through the correlated disturbances $v_{1},v_{2}$, we add $\color[rgb]{0,0,0}w_{2}$} to the predictor output. This leads to a predictor model $(w_{2},w_{3})\rightarrow(w_{1},w_{2})$, that satisfies the structural conditions for consistently estimating the target module. With this choice, $w_{i}=w_{2}$ and $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}=w_{3}$.
When analyzing the data-informativity conditions as developed in this paper, the vector signal $(w_{2},w_{3})$ would need appropriate excitation. Let us first consider possible locations $(x_{5},x_{3},x_{6})$ for unmeasured disturbance signals. Possible disturbance signals $x_{5}=e_{5}$ and $x_{6}=e_{6}$ will not provide excitation for $(w_{2},w_{3})$ as these disturbance signals have unmeasured paths to one of the output signals $w_{2}$ or $w_{1}$. Therefore their white noise sources are part of $e^{\!{\scriptscriptstyle\mathcal{Y}}}$, and cannot be used for excitation. A disturbance signal $x_{3}=e_{3}$ will lead to excitation for the input signal $w_{3}$.
@@
When applying the earlier results for full predictor model data-informativity, as presented in VandenHof&etal_IFAC:23 , the result would be that the present predictor model cannot satisfy the data-informativity conditions by adding a sufficient number of external signals only. Relying on the result of Proposition 5 in VandenHof&etal_IFAC:23 , we would need to extend the predictor model to $(w_{1},w_{2},w_{3},w_{5})\rightarrow(w_{1},w_{2})$, thus adding $w_{5}$ as input node signal to be measured, and also adding $w_{1}$ to the predictor input. This is based on the result of the mentioned Proposition, indicating that each loop in the network around nodes that appear both in the input and the output, should pass through an input node. Sufficient conditions for (full predictor model) data-informativity would then be that $w_{5}$ and $w_{3}$ are excited by either a noise signal $(e_{3},e_{5})$ or an excitation signals $(u_{3},u_{5})$, and that both $w_{1}$ and $w_{2}$ are excited by an external excitations signals $u_{1}$ and $u_{2}$. This illustrates how single module data-informativity can lead to significantly more relaxed conditions for instrumenting the network with sensors and actuators.
@@ 10 Conclusions
The components of $x$ that appear in the equation for $u_{k}$, $k\in\{i\cup\mathcal{D}^{c}\cup\mathcal{T}\}$ are then the $x$ components that are directly added to the concerned nodes in the original system representation, plus all those components $x$ that only pass through nodes in $u_{\!{\scriptscriptstyle\mathcal{F}}}$ before they reach $u_{k}$. This is a direct result of Gaussian elimination (immersion) of the nodes $u_{\!{\scriptscriptstyle\mathcal{F}}}$ (Dankers&etal_TAC:16, ).
@@
Since according to (15), $u_{\!{\scriptscriptstyle\mathcal{T}}}$ can be written as a function of $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ and $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$, substituting this in the above expression provides
@@
$\cdot\begin{bmatrix}u_{i}\\ u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\\ {\color[rgb]{0,0,0}x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}\end{bmatrix}$.
Denote $\nu:=\begin{bmatrix}u_{i}\\ u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\end{bmatrix}$. Then we can decompose $\nu$ according to $\nu={\color[rgb]{0,0,0}\nu_{x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}}+{\color[rgb]{0,0,0}\nu_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}}$ and using the fact that $\Delta_{j}^{b}(q,\theta)$ is parametrized independently from $\Delta G_{ji}$ and $\Delta_{j}^{a}$, we can rewrite $\Delta\bar{M}_{j}^{u}(q,\theta)\kappa_{u}^{[j]}$ as
@@
$\cdot\begin{bmatrix}{\color[rgb]{0,0,0}\nu_{x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}}+{\color[rgb]{0,0,0}\nu_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}}\\ {\color[rgb]{0,0,0}x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}\end{bmatrix}$
where $\Delta_{j}^{c}(q,\theta)$ replaces the term $\Delta_{j}^{a}(q,\theta)+\Delta_{j}^{b}(q,\theta)T_{s}(q)$. The expression for $\Delta\bar{M}_{j}^{u}(q,\theta)\kappa_{u}^{[j]}$ is now written as a summation of two terms that are mutually uncorrelated: one driven by $\color[rgb]{0,0,0}x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$} and $\color[rgb]{0,0,0}\nu_{x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}$} and one driven by its orthogonal complement. According to (33), for evaluating $\Delta\bar{M}_{j}(q,\theta)\kappa^{[j]}(t)$ a third term is added, given by $\Delta\bar{M}_{j}^{e}(q,\theta)\kappa_{e}^{[j]}$, leading to
@@
$\left(\begin{bmatrix}{\color[rgb]{0,0,0}v_{x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}}\\ {\color[rgb]{0,0,0}x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}\end{bmatrix}+\begin{bmatrix}{\color[rgb]{0,0,0}v_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}}\\ 0\end{bmatrix}\right)+\Delta\bar{M}_{j}^{e}(q,\theta)\kappa_{e}^{[j]}.$
Since both terms $\color[rgb]{0,0,0}v_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}$} and $\kappa_{e}^{[j]}$ are orthogonal to $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$, it follows now that $\bar{\mathbb{E}}\|\Delta\bar{M}_{j}(q,\theta)\kappa^{[j]}(t)\|^{2}=0$ implies $\begin{bmatrix}\Delta G_{ji}(q,\theta)&\Delta_{j}^{c}(q,\theta)&\Delta\bar{M}_{j}^{e}(q,\theta)\end{bmatrix}=0$ under the condition that $\Phi_{\mu^{[j]}}(\omega)\succ 0$ for almost all $\omega$ with $\mu^{[j]}$ given by
$\mu^{[j]}=\begin{bmatrix}{\color[rgb]{0,0,0}\nu_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}}\\ \kappa_{e}^{[j]}\end{bmatrix}$,
@@
As a consequence of the definition of $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ and the resulting $w_{\!{\scriptscriptstyle\mathcal{T}}}$, following Definition 6, the term $\Delta{\color[rgb]{0,0,0}\bar{M}_{j}^{w}}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}$ can be written as
$\Delta{\color[rgb]{0,0,0}\bar{M}_{j}^{w}}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}=\begin{bmatrix}\Delta\bar{G}_{ji}(q,\theta)&\Delta_{j}^{a}(q,\theta)&\Delta_{j}^{b}(q,\theta)\end{bmatrix}\begin{bmatrix}w_{i}\\ w_{\bar{\!}{\scriptscriptstyle\mathcal{D}}^{c}}\\ w_{\!{\scriptscriptstyle\mathcal{T}}}\end{bmatrix}$ (37)
with $w_{\bar{\!}{\scriptscriptstyle\mathcal{D}}^{c}}=w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\cap w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}$. By complementing $w_{\bar{\!}{\scriptscriptstyle\mathcal{D}}^{c}}$ with the remaining components of $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ and adding 0 entries to the accompanying $\Delta^{a}_{j}(q,\theta)$, this can be rewritten as
$\Delta{\color[rgb]{0,0,0}\bar{M}_{j}^{w}}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}=\begin{bmatrix}\Delta\bar{G}_{ji}(q,\theta)&\tilde{\Delta}_{j}^{a}(q,\theta)&\Delta_{j}^{b}(q,\theta)\end{bmatrix}\begin{bmatrix}w_{i}\\ w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\\ w_{\!{\scriptscriptstyle\mathcal{T}}}\end{bmatrix}$. (38)
Since according to (27), $w_{\!{\scriptscriptstyle\mathcal{T}}}$ can be written as a function of $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ and $\color[rgb]{0,0,0}x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$}, substituting this in the above expression provides